# CARE: Contrastive Alignment for ADL Recognition from Event-Triggered Sensor Streams

### v2

@@ Abstract
The recognition of Activities of Daily Living (ADLs) from event-triggered ambient sensors is an essential task in Ambient Assisted Living, yet existing methods remain constrained by representation-level limitations. Sequence-based approaches preserve temporal order of sensor activations but are sensitive to noise and lack spatial awareness, while image-based approaches capture global patterns and implicit spatial correlations but compress fine-grained temporal dynamics and distort sensor layouts. Naïve fusion (e.g., feature concatenation) fail to enforce alignment between sequence- and image-based representation views, underutilizing their complementary strengths. We propose Contrastive Alignment for ADL Recognition from Event-Triggered Sensor Streams (CARE), an end-to-end framework that jointly optimizes representation learning via Sequence–Image Contrastive Alignment (SICA) and classification via cross-entropy, ensuring both cross-representation alignment and task-specific discriminability. CARE integrates (i) time-aware, noise-resilient sequence encoding with (ii) spatially-informed and frequency-sensitive image representations, and employs (iii) a joint contrastive-classification objective for end-to-end learning of aligned and discriminative embeddings. Evaluated on three CASAS datasets, CARE achieves state-of-the-art performance (89.8% on Milan, 88.9% on Cairo, and 73.3% on Kyoto7) and demonstrates robustness to sensor malfunctions and layout variability, highlighting its potential for reliable ADL recognition in smart homes.
@@ I Introduction
Global increases in life expectancy are leading to aging societies, with a rising number of older adults who require continuous support from healthcare providers and their family members [30]. However, given the critical shortage of healthcare personnel, it is essential to support older adults in maintaining independence for as long as possible. [16]. An essential component of independent living is the ability to perform Activities of Daily Living (ADLs)—such as toileting, dressing, feeding, and cooking—that directly determine an individual’s quality of life. These functional abilities often decline with aging, and can be further deteriorated by aging-related chronic conditions [32].
Ambient Assisted Living (AAL) technologies have emerged to support ADL performance, encompassing systems for activity recognition, anomaly detection, and personalized prompting. Among these, ambient sensor-based systems are especially attractive for AAL, as they provide unobtrusive and privacy-preserving monitoring without the discomfort of wearables or the intrusiveness of cameras [15]. Embedded in home environments, ambient sensors offer passive, continuous, and cost-effective monitoring of motion, door usage, and environmental states, enabling scalable deployment in real-world homes [13, 12].
@@
*PSI: Physical Spatial Information, †NR: Noise Robust, ‡LO: Learning Objective, $\dagger$ SE: Sentence Encoder, BiL: BiLSTM, CE: Cross-Entropy.
@@
These limitations highlight a major gap: existing methods focus exclusively on either sequence or image encodings, treating temporal and spatial cues in isolation, and missing the opportunity to exploit their natural complementarity. Prior work has explored simple fusion strategies—such as feature concatenation or late fusion. While naïve fusion strategy (simple concatenation) could, in principle, combine both views, it ignores cross-view interactions and lacks explicit constraints to ensure that both representation views are effectively utilized, limiting its ability to fully leverage complementary information. To overcome this, we argue for a unified framework that can jointly leverage both perspectives within a shared latent space.
To this end, we propose CARE (A Contrastive Alignment Framework for ADL Recognition from Event-Triggered Sensor Streams), an end-to-end framework that unifies sequence- and image-based representations through supervised contrastive alignment and joint classification learning. At its core, CARE introduces SICA (Sequence–Image Contrastive Alignment), a supervised contrastive mechanism that aligns embeddings from the two representation views in a shared latent space. By enforcing both intra- and cross-view consistency, SICA produces richer, noise-resilient, and class-discriminative representations. By coupling alignment and classification within one optimization stage, CARE avoids the common two-step pipeline of representation pretraining followed by classifier training, thereby producing deployment-ready models that balance robustness, efficiency, and interpretability.
The main contributions of this work are threefold:
  • Robust temporal encoding. We design a time-aware and noise-resilient sequence encoding method that preserves global daily context while mitigating impact of spurious sensor firings, addressing the limitations of prior order-only encodings.
  • Spatially-informed image representation. We propose a dual-perspective image representation that integrates temporal dynamics with floorplan-based spatial cues and frequency-sensitive transitions, enabling richer and more interpretable activity patterns.
  • Cross-view contrastive alignment (SICA). We introduce a supervised contrastive alignment objective that unifies sequence- and image-based embeddings in a shared latent space, enforcing view-invariant and class-consistent representations.
  • Unified end-to-end framework (CARE). Building on SICA, we develop CARE, a one-stage framework that jointly optimizes representation learning and classification, eliminating multi-stage retraining and improving deployment readiness for real-world ADL systems.
@@ II Related Work
The integration of ambient sensors into assistive living technologies has shown strong potential to promote independent living and enable early detection of health deterioration [27]. Research has demonstrated that ambient sensor data, when analyzed with advanced machine learning techniques, can effectively classify and predict ADLs with high accuracy [1]. Ambient sensor-based systems are especially attractive for AAL because they are unobtrusive, cost-effective, and scalable, offering continuous monitoring of motion, door usage, and environmental states in naturalistic homes [13, 12]. Moreover, they align with broader trends toward privacy-preserving monitoring, respecting autonomy while enabling timely health alerts [17]. However, the event-triggered nature of such data makes ADL recognition inherently challenging: signals are sparse, irregular, and noisy, and older adults often perform the same activity with variations in order, speed, or completion [36]. These characteristics underscore the need for robust representation learning methods that can jointly capture fine-grained temporal dynamics and cross-sensor spatial context [21, 20, 2].
Table I summarizes representative encoding strategies of event-triggered sensor data for ADL recognition. Below, we review four main lines of work.
@@
Image-Based Encoding Approaches. An alternative line of work transforms event streams into images and applies convolutional encoders. BinaryCNN [35], DCNN [18] and GreyDCNN [28] map sensor activations into binary or grayscale activity images, , allowing CNNs to exploit global structures and long-range correlations of time series. These approaches are generally more robust to noise than sequence encoders, since image representations smooth over missing or spurious events. However, several limitations remain: (i) the commonly used 1D sensor indexing distorts the true 2D spatial layout, placing unrelated sensors adjacently and losing room-level semantics; (ii) temporal details are compressed into a single axis or pixel intensity, weakening fine-grained activity dynamics; and (iii) all sensor activations are treated equally, overlooking activation frequency as a discriminative cue.
@@
Other Input Representations. Graph-based approaches [31] represent sensor activations as nodes and transitions as edges, enabling explicit modeling of cross-sensor relationships. While effective on certain datasets, such models are computationally expensive due to the repeated graph construction and inference iterations, which limit their scalability in real-world deployments. More recently, textual encodings have emerged: Jüttner et al. [23] applied GPT-4 to infer ADLs and detect anomalies in CASAS [12], while TDOST [34] tokenizes event sequences into words and leverages a pretrained sentence transformer followed by a BiLSTM and classifier. Notably, TDOST achieves state-of-the-art performance across several benchmarks without explicit temporal features, underscoring the potential of language-inspired representations. However, when this method plugin the time information into the words, the performance is not improved. This highlights an open question: how to exploit temporal information in non-traditional encodings like text while preserving underlying activity patterns.
Contrastive Learning for ADL Recognition. Contrastive learning was initially designed to transform features so that similar samples are positioned close to each other in a transformed feature space [9]. While it has shown promise in domains such as vision and NLP [8, 24, 5], the adoption in ambient sensing remains limited [6, 7]. Existing ADL methods typically focus on single representation view or naïve fusion, where features from different views are concatenated and trained with cross-entropy loss. Such strategies lack explicit constraints to ensure that both views are effectively leveraged, and thus miss the opportunity to exploit cross-view complementarity. In contrast, our proposed method, the CARE framework, introduces a supervised contrastive alignment objective (SICA) that explicitly enforces consistency between sequence- and image-based embeddings in a shared latent space, addressing the limitations of prior methods and enabling more robust and discriminative ADL representations.
@@ III-A Time-Aware and Noise-Resilient Sequence Encoding
Temporal Binning. A critical limitation of prior sequence encodings is their inability to handle temporal information. Existing methods[19, 22, 28, 31, 34] typically: (i) ignore timestamps entirely, discarding daily context; (ii) use inter-event intervals $\Delta t_{i}=t_{i}-t_{i-1}$, which capture local timing gaps but fail to reflect absolute daily temporal positioning; or (iii) combine inter-event intervals with activity begin times $(t_{\mathrm{start}},\Delta t_{i})$, which anchors the sequence to its onset and preserves absolute time-of-day information at the activity level, but does not explicitly associate each sensor event with its own global timestamp. In contrast, our method assigns a coarse-grained temporal embedding to every event. For each raw time-of-day $t_{i}\in\mathbb{R}^{+}$, we compute
@@
Figure 2: Frequency-based Event Filtering on Cairo
@@
Sequence Representation. To ensure consistency across heterogeneous sensor types, binary states (“ON/OPEN/PRESENT”) are mapped to 1 and others to 0, while continuous values such as temperature are normalized to $[0,1]$. Each activity segment is further padded or truncated to a fixed length $L$ to enable batch processing. Each event is then represented by concatenating a one-hot encoding of its sensor ID, the temporal index $\tau_{i}$, and the normalized signal state. The resulting enriched sequence is processed by a sequence encoder (e.g., LSTM) to generate temporal embeddings $h_{t}$ that jointly capture sensor identity and contextualized temporal dependencies (Fig. 3).
@@ III-B Spatially-Informed and Frequency-Sensitive Image Representation
Figure 3: Sequence Representation.
Prior studies [18, 28] typically represent activities as binary or grayscale matrices, with the $x$-axis denoting event order and the $y$-axis denoting sensor IDs. This design has several advantages: it preserves the temporal order of events and, to some extent, encodes implicit spatial relations through the vertical sensor index. However, the spatial information is inaccurate, as the one-dimensional $y$-axis does not reflect the true two-dimensional floorplan. Sensors located in different rooms may be placed adjacently in the $y$-axis, leading to semantic confusion. Furthermore, such encodings treat all sensor activations equally and ignore their activation frequencies, failing to distinguish frequent transitions from sporadic noise.
@@
Temporal Image. For an activity segment with events $\{e_{i}\}_{i=1}^{n}$, where $e_{i}=(\tau_{i},s_{i},a_{i})$ denotes temporal index, sensor ID, and normalized activation signal, we construct a temporal image $I^{\mathrm{temp}}$ such that $I^{\mathrm{temp}}(x_{i},y_{i})=\phi(\mathrm{idx}(\tau_{i}),s_{i},a_{i}),$ where $x_{i}$ corresponds to the event order ($x_{1}$ denotes the first timestamp, followed sequentially up to $x_{L}$ after padding or truncation), $y_{i}$ is the index of sensor $s_{i}$, and $\mathrm{idx}(\tau_{i})$ is the discretized time-bin index of timestamp. The mapping function $\phi(\cdot)$ integrates the time-bin index, sensor identity, and signal state into color channels (e.g., motion = blue, door = yellow, temperature = red). This representation highlights temporal dynamics while differentiating sensor modalities.
Spatial Image. Each sensor $s_{j}$ is associated with normalized floorplan coordinates $(u_{j},v_{j})\in[0,256)^{2}$. We construct a spatial image $I^{\mathrm{spat}}$ where nodes represent sensor positions and edges represent transitions between consecutive events. Nodes are rendered according to their activation frequency $f_{j}$ (already defined in Sec. III-A), normalized by $\max_{k}f_{k}$, with darker colors indicating more frequently activated sensors. For edges, the transition from $s_{j}$ to $s_{k}$ is assigned weight
@@
where $c_{jk}$ is the number of observed transitions between $s_{j}$ and $s_{k}$ in the segment, with darker edges denoting more frequent transitions. States such as “OFF”, “CLOSE”, and “ABSENT” are excluded to reduce clutter. This representation leverages the floorplan to embed spatial context and highlights recurrent patterns while suppressing sporadic noise.
Unified Image Representation. The temporal and spatial images are concatenated to form a composite image $I^{\mathrm{comp}}=\text{concat}(I^{\mathrm{temp}},I^{\mathrm{spat}}),$ which jointly encodes temporal ordering, spatial configuration, and frequency awareness. The composite image is then processed by a CNN encoder to produce an embedding that is spatially-informed, frequency-sensitive, and robust to noise (Fig. 4).

### v3

@@ Abstract
The recognition of Activities of Daily Living (ADLs) from event-triggered ambient sensors is an essential task in Ambient Assisted Living, yet existing methods remain constrained by representation-level limitations. Sequence-based approaches preserve temporal order of sensor activations but are sensitive to noise and lack spatial awareness, while image-based approaches capture global patterns and implicit spatial correlations but compress fine-grained temporal dynamics and distort sensor layouts. Naïve fusion (e.g., feature concatenation) fails to enforce alignment between sequence- and image-based representation views, underutilizing their complementary strengths. We propose Contrastive Alignment for ADL Recognition from Event-Triggered Sensor Streams (CARE), an end-to-end framework that jointly optimizes representation learning via Sequence–Image Contrastive Alignment (SICA) and classification via cross-entropy, ensuring both cross-representation alignment and task-specific discriminability. CARE integrates (i) time-aware, noise-resilient sequence encoding with (ii) spatially-informed and frequency-sensitive image representations, and employs (iii) a joint contrastive-classification objective for end-to-end learning of aligned and discriminative embeddings. Evaluated on three CASAS datasets, CARE achieves state-of-the-art performance (89.8% on Milan, 88.9% on Cairo, and 73.3% on Kyoto7) and demonstrates robustness to sensor malfunctions and layout variability, highlighting its potential for reliable ADL recognition in smart homes. We release our code at https://github.com/Jhziiiig/CARE.
@@ I Introduction
Global increases in life expectancy are leading to aging societies, with a rising number of older adults who require continuous support from healthcare providers and their family members [28]. However, given the critical shortage of healthcare personnel, it is essential to support older adults in maintaining independence for as long as possible [12]. An essential component of independent living is the ability to perform Activities of Daily Living (ADLs)—such as toileting, dressing, feeding, and cooking—that directly affect an individual’s quality of life. These functional abilities often decline with aging and can be further compromised by age-related chronic conditions [30].
Ambient Assisted Living (AAL) technologies have emerged to support ADL performance, encompassing systems for activity recognition, anomaly detection, and personalized prompting. Among these, ambient sensor-based systems are especially attractive for AAL, as they provide unobtrusive and privacy-friendly monitoring without wearables’ discomfort or cameras’ intrusiveness [11]. Embedded in home environments, ambient sensors offer passive, continuous, and cost-effective monitoring of motion, door usage, and environmental states, enabling scalable deployment in real-world homes  [9].
@@
*PSI: Physical Spatial Information, †NR: Noise Robust, ‡LO: Learning Objective, $\star$ SE: Sentence Encoder, BiL: BiLSTM, CE: Cross-Entropy.
@@
These limitations highlight a major gap: existing methods focus exclusively on either sequence or image encodings, treating temporal and spatial cues in isolation, and missing the opportunity to exploit their natural complementarity. Prior work has explored simple fusion strategies—such as feature concatenation or late fusion. While naïve fusion strategy (simple concatenation) could, in principle, combine both views, it ignores cross-view interactions and lacks explicit constraints to ensure that both representation views are effectively utilized, limiting its ability to fully leverage complementary information. To overcome this, we argue for a unified framework that can jointly leverage both sequence and image perspectives within a shared latent space.
To this end, we propose CARE (A Contrastive Alignment Framework for ADL Recognition from Event-Triggered Sensor Streams), an end-to-end framework that unifies sequence- and image-based representations through supervised contrastive alignment and joint classification learning. CARE is novel at the framework level: it treats sequence and image as two complementary but heterogeneous views of the same event stream, and introduces an explicit cross-view alignment principle (SICA: Sequence–Image Contrastive Alignment) to learn a shared latent space that is simultaneously view-consistent and class-discriminative, jointly optimized with task supervision in one stage.
The main contributions of this work are:
  • A cross-view alignment framework for event-triggered ADL recognition (CARE). We propose a backbone-agnostic framework that aligns sequence and image views of the same event stream, and jointly optimizes cross-view alignment and task supervision in one stage.
  • Sequence–Image Contrastive Alignment (SICA). We introduce a supervised contrastive alignment objective that unifies sequence- and image-based embeddings in a shared latent space, enforcing view-invariant and class-consistent representations.
  • Robust temporal encoding. We design a time-aware, noise-resilient sequence encoding that preserves global daily context while suppressing spurious sensor firings, addressing the limitations of prior order-only encodings.
  • Spatially-informed and frequency-sensitive image representation. We propose a dual-perspective image representation that integrates temporal patterns with sensor-layout spatial cues and frequency-sensitive transitions, yielding richer and more interpretable activity signatures.
@@ II Related Work
The integration of ambient sensors into assistive living technologies has shown strong potential to promote independent living and enable early detection of health deterioration [25]. Research has demonstrated that ambient sensor data, when analyzed with advanced machine learning techniques, can effectively classify and predict ADLs with high accuracy [1]. Ambient sensor-based systems are especially attractive for AAL because they are unobtrusive, cost-effective, and scalable, offering continuous monitoring of motion, door usage, and environmental states in naturalistic homes [9]. Moreover, they align with broader trends toward privacy-preserving monitoring, respecting autonomy while enabling timely health alerts [15]. However, the event-triggered nature of such data makes ADL recognition inherently challenging: signals are sparse, irregular, and noisy, and older adults often perform the same activity with variations in order, speed, or completion [34]. These characteristics underscore the need for robust representation learning methods that can jointly capture fine-grained temporal dynamics and cross-sensor spatial context [20, 19, 2]. Table I summarizes representative encoding strategies of event-triggered sensor data for ADL recognition. Below, we review four main lines of work.
@@
Image-Based Encoding Approaches. An alternative line of work transforms event streams into images and applies convolutional encoders. BinaryCNN [33], DCNN [16] and GreyDCNN [26] map sensor activations into binary or grayscale activity images, allowing CNNs to exploit global structures and long-range correlations of time series. These approaches are generally more robust to noise than sequence encoders, since image representations smooth over missing or spurious events. However, several limitations remain: (i) the commonly used 1D sensor indexing distorts the true 2D spatial layout, placing unrelated sensors adjacently and losing room-level semantics; (ii) temporal details are compressed into a single axis or pixel intensity, weakening fine-grained activity dynamics; and (iii) all sensor activations are treated equally, overlooking activation frequency as a discriminative cue.
@@
Other Input Representations. Graph-based approaches [29] represent sensor activations as nodes and transitions as edges, enabling explicit modeling of cross-sensor relationships. While effective on certain datasets, such models are computationally expensive due to the repeated graph construction and inference iterations, which limit their scalability in real-world deployments. More recently, textual encodings have emerged: Jüttner et al. [JüttnerFleigBuchmann+2025+173+187] applied GPT-4 to infer ADLs and detect anomalies in CASAS [9], while TDOST [32] tokenizes event sequences into words and leverages a pretrained sentence transformer followed by a BiLSTM and classifier. Notably, TDOST achieves state-of-the-art performance across several benchmarks without explicit temporal features, underscoring the potential of language-inspired representations. However, when this method plugs the time information into the words, the performance is not improved. This highlights an open question: how to exploit temporal information in non-traditional encodings like text while preserving underlying activity patterns.
Contrastive Learning for ADL Recognition. Contrastive learning learns representations by bringing semantically similar samples closer and pushing dissimilar ones apart in the embedding space [7]. It has achieved strong results in vision and NLP [21], and has increasingly been explored for human activity recognition (HAR). Representative examples include CPC [18], which applies contrastive objectives to learn temporal features, and multi-view HAR frameworks such as MuJo [13] and SEAL [14], which align representations across modalities or between sensor signals and label semantics. However, these studies primarily target continuously sampled wearable or rich multimodal streams (e.g., IMU, video, pose, text), and do not directly address event-triggered ambient sensor streams that are sparse, irregular, and noisy. As a result, the adoption of contrastive learning for ambient, event-triggered ADL recognition remains limited [5, 6], where most prior methods still adopt a single-view encoding (sequence-only or image-only), or use naïve fusion (feature concatenation with cross-entropy) when multiple views are available. Such designs provide no explicit mechanism to align heterogeneous views, and may underutilize complementary temporal and spatial cues. CARE fills this gap by introducing a supervised cross-view contrastive objective to align sequence and image embeddings in a shared latent space.
@@ III-A Time-Aware and Noise-Resilient Sequence Encoding
Temporal Binning. Prior sequence encodings often struggle to handle temporal information. Existing methods[17, 23, 26, 29, 32] typically: (i) ignore timestamps entirely, discarding daily context; (ii) use inter-event intervals $\Delta t_{i}=t_{i}-t_{i-1}$, which capture local timing gaps but fail to reflect absolute daily temporal positioning; or (iii) combine inter-event intervals with activity begin times $(t_{\mathrm{start}},\Delta t_{i})$, which anchors the sequence to its onset and preserves absolute time-of-day information at the activity level, but does not explicitly associate each sensor event with its own global timestamp. In contrast, our method assigns a coarse-grained temporal embedding to every event. For each raw time-of-day $t_{i}\in\mathbb{R}^{+}$, we compute
@@
Figure 2: Frequency-based Event Filtering on Cairo.
@@
Sequence Representation. To ensure consistency across heterogeneous sensor types, binary states (e.g., “ON”, “OPEN”, “PRESENT”) are mapped to 1 and others to 0, while continuous values such as temperature (e.g., 20.5 and 19.5) are normalized to $[0,1]$. Each activity segment is further padded or truncated to a fixed length $L$ to enable batch processing. Each event is then represented by concatenating a one-hot encoding of its sensor ID, the temporal index $\tau_{i}$, and the normalized signal state. The resulting enriched sequence is processed by a sequence encoder (e.g., LSTM) to generate temporal embeddings $h_{t}$ that jointly capture sensor identity and contextualized temporal dependencies (Fig. 3).
Figure 3: Sequence Representation.
@@ III-B Spatially-Informed and Frequency-Sensitive Image Representation
Prior studies [16, 26] typically represent activities as binary or grayscale matrices, with the $x$-axis denoting event order and the $y$-axis denoting sensor IDs. This design has several advantages: it preserves the temporal order of events and, to some extent, encodes implicit spatial relations through the vertical sensor index. However, the spatial information is inaccurate, as the one-dimensional $y$-axis does not reflect the true two-dimensional floor-plan. Sensors located in different rooms may be placed adjacently in the $y$-axis, leading to semantic confusion. Furthermore, such encodings treat all sensor activations equally and ignore their activation frequencies, failing to distinguish frequent transitions from sporadic noise.
@@
Temporal Image. For an activity segment with events $\{e_{i}\}_{i=1}^{n}$, where $e_{i}=(\tau_{i},s_{i},a_{i})$ denotes temporal index, sensor ID, and normalized activation signal, we construct a temporal image $I^{\mathrm{temp}}$ such that $I^{\mathrm{temp}}(x_{i},y_{i})=\phi(\mathrm{idx}(\tau_{i}),s_{i},a_{i}),$ where $x_{i}$ corresponds to the event order ($x_{1}$ denotes the first timestamp, followed sequentially up to $x_{L}$ after padding or truncation), $y_{i}$ is the index of sensor $s_{i}$, and $\mathrm{idx}(\tau_{i})$ is the discretized time index of timestamp. The mapping function $\phi(\cdot)$ integrates the time index, sensor identity, and signal state into color channels (e.g., motion = blue, door = yellow, temperature = red). This representation highlights temporal dynamics while differentiating sensor modalities.
Spatial Image. Each sensor $s_{j}$ is associated with normalized floor-plan coordinates $(u_{j},v_{j})\in[0,256)^{2}$. We construct a spatial image $I^{\mathrm{spat}}$ where nodes represent sensor positions and edges represent transitions between consecutive events. Nodes are rendered according to their activation frequency $f_{j}$ (already defined in Sec. III-A), normalized by $\max_{k}f_{k}$, with darker colors indicating more frequently activated sensors. For edges, the transition from $s_{j}$ to $s_{k}$ is assigned weight
@@
where $c_{jk}$ is the number of observed transitions between $s_{j}$ and $s_{k}$ in the segment, with darker edges denoting more frequent transitions. States such as “OFF”, “CLOSE”, and “ABSENT” are excluded to reduce clutter. This representation leverages the floor-plan to embed spatial context and highlights recurrent patterns while suppressing sporadic noise.
Unified Image Representation. The temporal and spatial images are concatenated to form a composite image $I^{\mathrm{comp}}=\text{concat}(I^{\mathrm{temp}},I^{\mathrm{spat}}),$ which jointly encodes temporal ordering, spatial configuration, and frequency awareness. The composite image is then processed by a CNN encoder to produce an embedding that is spatially-informed, frequency-sensitive, and robust to noise (Fig. 4). The floor-plan background is shown for illustration only and is not included in spatial images.